OpenStudy (anonymous):

let f (x) = 4/x^2 + 5 ... g (x) = Sqrt x find f o g and g o f.

OpenStudy (anonymous):

fog = f((g)) = 4x + 5 gof = g((f)) = Sqrt(4x^2+5)

OpenStudy (anonymous):

where is sqrt gone in function f o g ?

OpenStudy (anonymous):

hmm 'cux (Sqrtx)^2 = x right??

OpenStudy (anonymous):

You mean we canceled out sqrt with ^2 ? and please also explain the procedure how we can do it.

OpenStudy (anonymous):

yep fog mean that you put g(x) in stead of x in f(x) so x in f(x) should change into Sqrtx get it?

OpenStudy (anonymous):

and also where is division sign has gone ?

OpenStudy (anonymous):

oh nooo it's not gone I just saw it wrong lol

OpenStudy (anonymous):

it should be 4/x + 5 then

OpenStudy (anonymous):

yea, thnx and 2nd one?

OpenStudy (anonymous):

Sqrt(4/x^2+5)

OpenStudy (anonymous):

in 2nd one every value is included in sqrt , right?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

\[fog(x)=\frac4x+5\] \[gof(x)=\sqrt{\frac{4}{x^2}+5}\]

OpenStudy (anonymous):

thnx cinar